a. When a best-selling book was first released in paperback, the Hercules Bookstore chain seized a profit opportunity by setting a selling price of $9 per book (well above Hercules’ $5 average cost per book). With paperback demand given by P = 15 - 5Q, the chain enjoyed sales of Q = 12 thousand books per week. (Note: Q is measured in thousands of books.) Draw the demand curve and compute the bookstore’s profit and the total consumer surplus.
b. For the first time, Hercules has begun selling books online—in response to competition from other online sellers and in its quest for new profit sources. The average cost per book sold online is only $4. As part of its online selling strategy, it sends weekly e-mails to preferred customers announcing which books are new in paperback. For this segment, it sets an average price (including shipping) of $12. According to the demand curve in part (a), only the highest value consumers (whose willingness to pay is $12 or more) purchase at this price. Check that these are the first 6 thousand book buyers on the demand curve. In turn, because of increased competition, Hercules has reduced its store price to $7 per book. At P = $7, how many books are bought in Hercules’ stores? (Make sure to exclude online buyers from your demand curve calculation.) Compute Hercules’ total profit. Then compute the sum of consumer surplus from online and in-store sales. Relative to part (a), has the emergence of online commerce improved the welfare of book buyers as a whole? Explain.
Answer #1

Answer to Question a:

Given details are:

Price of one book = $9

Cost of one book= $5

Demand for the book is P=15-5Q

Computation:

Demand for books is P=15-5Q substituting value of Q in the equation:

P=15-5(12)

p=15-60=45

Hence the total demand for books is 45000

Total profit earned by the firm is as follows:

Profit =45000*($9-$5)

Profit= $180,000.

Computation of consumer surplus:

Consumer surplus is nothing but the willingness of the consumer to pay in excess of the market price for a commodity. in the given case one book can be sold for $9 and the total demand for book is 45000. The consumer surplus is the also equal to the profit earned by the firm or it is $4 (i.e.$9-$5) per book. since the price of the book is to changing as the quantity demanded increases. in simple terms consumer surplus is the excess of price over the average cost of producing a product.